Calculate Reynolds number, if a fluid having viscosity of 0.4 Ns/m2 and relative density of 900 kg/m3 through a pipe of 20mm with a velocity of 2.5 m.
1
Expert's answer
2020-04-19T07:40:03-0400

Reynold number is used to determine the flow of fluid which is laminar or turbulent


"R_e= \frac{\rho vd}{\mu}" , here "\rho= density" of fluid="900 \frac{kg}{m^3}" , v= velocity of the fluid =2.5 m/s


d= diameter of pipe=20 mm= 0.02m, "\mu" = coefficient of viscosity= 0.4 "\frac {Ns}{m^2}"


"R_e= \frac{\rho vd}{\mu}"


"R_e= \frac{900\times 2.5\times 0.02}{0.4}"


"R_e=" 112.5

So, The Reynold number for this fluid flow is =112.5